OpenStudy (anonymous):

How do I use elimination, and substitution? I get it, somewhat, But i need to scrub up on it!

OpenStudy (anonymous):

Elimination: Ax+B=C Dx-B=F __________ now add the equations to get (Ax+Dx)=(C++F) because the +/- cancels itself.

OpenStudy (anonymous):

substitution means you "substute the value of one variable from another equaiton

OpenStudy (katrinakaif):

Substiution is when you plug in something - we need to equations for it. For example, y + x = 530 and 5(x) + 3.75(y) = 2412.50. Then we can substitute 530 -x = y for why..leaving us to: 5x + 3.75( 530-x) = 2412.50. Thats Substitution

OpenStudy (anonymous):

Substitution: Ax+By=C Dx=y ________ substitute y into the first equation to get Ax+B(Dx)=C

OpenStudy (anonymous):

elimination is usually done when you can see that the differnce between 2 coefficients is a small factor (mult)..whereas subsitution is the more general methodi.e. works for non linear cases as well

OpenStudy (katrinakaif):

Elimination is when you have 2 equations that can cancel one of the variables. Like - y + 2x = 6 + y + 7x = 13 w you can see that - y from the top expression can be cancelled out with the bottom postive y. then we add the x terms and constant terms to give us one variable. then we can use it to solve for the other variable as for Y by pluging in for ONE of the eany equation